Session: Mar 2010 (session 1)
Wavelength: 18 cm


To give feedback on this experiment, click here

Experiment Feedback

Medicina (Roda at Sun Mar 07 16:05:18 2010) Success:

Westerbork (Geert Kuper at Sun Mar 07 16:38:46 2010) Success: RFI
Only one RT in the adder (RT7) as requested by PI. RFI at 1610.49 MHz. No signal from 11:04 UT til 12:14 UT due to backend problems. After 12:04 til end of this project no signal anymore in our subband 0 which is linked to BBC1 lower sideband and BBC2 lower sideband.

Effelsberg (dg at Mon Mar 08 10:05:08 2010) Minor failures: RECO LATE
Started late 13:36, mk5 server had crashed. Unusually large number of formatter comm problems, should not affect data

Torun (Kaz Borkowski at Mon Mar 08 13:19:26 2010) Success:
No known problems.

Onsala (Michael Lindqvist at Mon Mar 08 16:33:42 2010) Minor failures: RECO
Lost some time at start due to Mark5. Could not append to old disk. 2010.066.10:08:20.06?ERROR m5 -900 Can't append recording. A lot of RFI.

Noto (P. Cassaro at Fri Mar 26 11:56:59 2010) Minor failures: TSYS
The bbc #8 have no Tsys measures.

Jodrell_Bank (alastair gunn at Mon Mar 29 10:57:37 2010) Minor failures: ANT
Frequent telescope encoder faults giving brief periods off source. Telescope parked briefly at 11:00 UT to try and clear encoder faults. No other known problems.